OpenStudy (anonymous):

use shell method to find the volume of the following solid. The ellipsoid formed when the ellipse x^2+11y^2=121 is revolved about the y-axis. Please show steps on how problem is done, thanks!

OpenStudy (anonymous):

to simplify things, we can just take the volume of the region in the first quadrant, rotate it about the y-axis then multiply by 2... do you agree?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

yes go ahead!

OpenStudy (anonymous):

\[4\pi \int\limits_{0}^{11}x \sqrt{11-x^2/11}dx\]

OpenStudy (anonymous):

it's actually v=\[484\sqrt{11} \pi/3\]